OpenStudy (anonymous):

how much simple interest is earned in one year from two accounts bearing 4% and 6% respectively? The first contains 3000 and the second 2500

OpenStudy (wolf1728):

3,000 + 3,000 * .04 = 3,120 2,500 + 2,500 * .06 = 2,650

OpenStudy (anonymous):

so i just a them 3120+2650=5770

OpenStudy (preetha):

Yes.

OpenStudy (preetha):

Actually, the interest is just 3000*.04 + 2500*06.